# Irida-Graphene Phonon Thermal via Non-equilibrium Molecular Dynamics Simulations

### v1

@@ 3.3 Vibrational Spectrum
Figure 5: Calculated vibrational density of states (VDOS) for pristine graphene (gray baseline) and Irida-G (black line). Panels (a), (b), and (c) illustrate the contributions from longitudinal (LA/LO), transverse (TA/TO), and flexural (ZA/ZO) modes, respectively. Panel (d) shows the total VDOS.
Flexural phonons make the most important contributions to the VDOS in the low-frequency regime (below 10 THz), likely serving as the primary heat carriers in Irida-G. With their longer wavelengths, low-frequency phonons typically play a crucial role in heat transport in nanomaterials. These phonons undergo fewer scatterings and travel longer distances within the material, carrying significant energy and contributing more effectively to thermal transport. Moreover, the importance of flexural polarization (ZA) phonons, due to out-of-plane oscillations, in thermal transport has been well-documented for graphene [48, 73, 74, 38], and other 2D materials [51].
@@ 3.4 Phonon Group Velocity
where $\omega$ represents the frequency of a given mode and $q$ denotes the wave vector. Figure 6 illustrates the absolute values of the phonon group velocities as a function of frequency for Irida-G (black circles) and graphene (gray squares), both with the same number of atoms ($48$) and similar dimensions. Notably, data points between $\sim 2\times 10^{3}$ and $3\times 10^{3}$ m/s, ranging up to 15 THz for Irida-G and up to 35 THz for graphene, correspond to the acoustic modes, which are the primary contributors to thermal transport, as previously discussed. The data indicate that the phonon group velocities of Irida-G are lower than those of graphene. The average group velocities are approximately 410 m/s and 780 m/s for Irida-G and graphene, respectively. Since lattice thermal conductivity is proportional to the square of the group velocities [75], this highlights the significant difference in thermal conductivity between these two materials. These results are consistent with the observed lower thermal conductivity of Irida-G relative to graphene.
Figure 6: Phonon group velocities as a function of frequency calculated from phonon dispersion relations using harmonic approximation.
@@ 4 Conclusions
The authors declare no conflict of interest.
@@ Acknowledgements
This work received partial support from Brazilian agencies CAPES, CNPq, and FAPDF. I.M.F. thanks the National Council for Scientific and Technological Development (CNPq) for grant no. 153604/2024-7. L.D.M. acknowledges the support of the High Performance Computing Center at UFRN (NPAD/UFRN). D. S. G. acknowledges the Center for Computing in Engineering and Sciences at Unicamp for financial support through the FAPESP/CEPID Grant #2013/08293-7. L.A.R.J. L.A.R.J. acknowledges the financial support from FAP-DF $00193.00001808/2022-71$ and $00193-00001857/2023-95$ grant and FAPDF-PRONEM grant $00193.00001247/2021-20$, and CNPq grant $350176/2022-1$. M.L.P.J. acknowledges the financial support of the FAP-DF grant 00193-00001807/2023-16. Thanks also to CENAPAD-SP (National High-Performance Center in São Paulo, State University of Campinas – UNICAMP, project: proj960) and NACAD (High-Performance Computing Center, Lobo Carneiro Supercomputer, Federal University of Rio de Janeiro – UFRJ, project: a22002) for the computational support provided.

### v2

@@ 3.3 Vibrational Spectrum
Fig.  5: Calculated vibrational density of states (VDOS) for pristine graphene (gray baseline) and Irida-G (black line). Panels (a), (b), and (c) illustrate the contributions from longitudinal (LA/LO), transverse (TA/TO), and flexural (ZA/ZO) modes, respectively. Panel (d) shows the total VDOS.
Flexural phonons make the most important contributions to the VDOS in the low-frequency regime (below 10 THz), likely serving as the primary heat carriers in Irida-G. With their longer wavelengths, low-frequency phonons typically play a crucial role in heat transport in nanomaterials. These phonons undergo fewer scatterings and travel longer distances within the material, carrying significant energy and contributing more effectively to thermal transport. Moreover, the importance of flexural polarization (ZA) phonons, due to out-of-plane oscillations, in thermal transport has been well-documented for graphene 48, 73, 74, 38, and other 2D materials 51.
@@ 3.4 Phonon Group Velocity
where $\omega$ represents the frequency of a given mode and $q$ denotes the wave vector. Figure 6 illustrates the absolute values of the phonon group velocities as a function of frequency for Irida-G (black circles) and graphene (gray squares), both with the same number of atoms ($48$) and similar dimensions. Notably, data points between $\sim 2\times 10^{3}$ and $3\times 10^{3}$ m/s, ranging up to 15 THz for Irida-G and up to 35 THz for graphene, correspond to the acoustic modes, which are the primary contributors to thermal transport, as previously discussed. The data indicate that the phonon group velocities of Irida-G are lower than those of graphene. The average group velocities are approximately 410 m/s and 780 m/s for Irida-G and graphene, respectively. Since lattice thermal conductivity is proportional to the square of the group velocities 75, this highlights the significant difference in thermal conductivity between these two materials. These results are consistent with the observed lower thermal conductivity of Irida-G relative to graphene.
Fig.  6: Phonon group velocities as a function of frequency calculated from phonon dispersion relations using harmonic approximation.
@@ 4 Conclusions
There are no conflicts to declare.
Data availability
The data that support the findings of this study are available from the corresponding author, M.L.P.J., upon reasonable request
@@ Acknowledgements
This work received partial support from Brazilian agencies CAPES, CNPq, and FAPDF. I.M.F. thanks the National Council for Scientific and Technological Development (CNPq) for grant no. 153604/2024-7. R.M.T. acknowledges the support of the MackGraphe. L.D.M. acknowledges the support of the High Performance Computing Center at UFRN (NPAD/UFRN). D. S. G. acknowledges the Center for Computing in Engineering and Sciences at Unicamp for financial support through the FAPESP/CEPID Grant #2013/08293-7. L.A.R.J. acknowledges the financial support from FAP-DF 00193.00001808/2022-71 and 00193-00001857/2023-95 grant and FAPDF-PRONEM grant 00193.00001247/2021-20, and CNPq grant 350176/2022-1. M.L.P.J. acknowledges the financial support of the FAP-DF grant 00193-00001807/2023-16. Thanks also to CENAPAD-SP (National High-Performance Center in São Paulo, State University of Campinas – UNICAMP, project: proj960) and NACAD (High-Performance Computing Center, Lobo Carneiro Supercomputer, Federal University of Rio de Janeiro – UFRJ, project: a22002) for the computational support provided.
Notes and references
  • Radhakrishnan et al. 2024 S. Radhakrishnan, P. P. Das, A. Alam, S. P. Dwivedi and V. Chaudhary, Proc. Inst. Mech. Eng. Pt. C J. Mechan. Eng. Sci., 2024, 09544062241245018.
  • Duan et al. 2023 F. Duan, D. Wei, A. Chen, X. Zheng, H. Wang and G. Qin, Nanoscale, 2023, 15, 1459–1483.
  • Huang et al. 2020 P. Huang, P. Zhang, S. Xu, H. Wang, X. Zhang and H. Zhang, Nanoscale, 2020, 12, 2309–2327.
  • Liu and Fan 2023 F. Liu and Z. Fan, Chem. Soc. Rev., 2023, 52, 1723–1772.
  • Salahdin et al. 2022 O. D. Salahdin, H. Sayadi, R. Solanki, R. M. R. Parra, M. Al-Thamir, A. T. Jalil, S. E. Izzat, A. T. Hammid, L. A. B. Arenas and E. Kianfar, Appl. Phys. A, 2022, 128, 703.
  • Tiwari et al. 2020 S. K. Tiwari, S. Sahoo, N. Wang and A. Huczko, J. Sci.: Adv. Mater. Devices, 2020, 5, 10–29.
  • Novoselov et al. 2004 K. S. Novoselov, A. K. Geim, S. V. Morozov, D.-e. Jiang, Y. Zhang, S. V. Dubonos, I. V. Grigorieva and A. A. Firsov, Science, 2004, 306, 666–669.
  • Mbayachi et al. 2021 V. B. Mbayachi, E. Ndayiragije, T. Sammani, S. Taj, E. R. Mbuta et al., Results Chem., 2021, 3, 100163.
  • Ogawa et al. 2023 S. Ogawa, S. Fukushima and M. Shimatani, Materials, 2023, 16, 2005.
  • Zhu et al. 2021 Q. Zhu, Z. Wang, H. Zeng, T. Yang and X. Wang, Compos. Part A: Appl. Sci. Manuf., 2021, 142, 106240.
  • Egbedina et al. 2022 A. O. Egbedina, O. P. Bolade, U. Ewuzie and E. C. Lima, J. Environ. Chem. Eng., 2022, 10, 107260.
  • Speranza 2021 G. Speranza, Nanomaterials, 2021, 11, 967.
  • Kanungo et al. 2022 S. Kanungo, G. Ahmad, P. Sahatiya, A. Mukhopadhyay and S. Chattopadhyay, npj 2D Mater. Appl., 2022, 6, 83.
  • Wang et al. 2015 Z. Wang, X.-F. Zhou, X. Zhang, Q. Zhu, H. Dong, M. Zhao and A. R. Oganov, Nano Lett., 2015, 15, 6182–6186.
  • Fan et al. 2019 Q. Fan, D. Martin-Jimenez, D. Ebeling, C. K. Krug, L. Brechmann, C. Kohlmeyer, G. Hilt, W. Hieringer, A. Schirmeisen and J. M. Gottfried, J. Am. Chem. Soc., 2019, 141, 17713–17720.
  • Baughman et al. 1987 R. Baughman, H. Eckhardt and M. Kertesz, J. Chem. Phys., 1987, 87, 6687–6699.
  • Hu et al. 2022 Y. Hu, C. Wu, Q. Pan, Y. Jin, R. Lyu, V. Martinez, S. Huang, J. Wu, L. J. Wayment, N. A. Clark et al., Nat. Synth., 2022, 1, 449–454.
  • Hudspeth et al. 2010 M. A. Hudspeth, B. W. Whitman, V. Barone and J. E. Peralta, ACS Nano, 2010, 4, 4565–4570.
  • Fan et al. 2021 Q. Fan, L. Yan, M. W. Tripp, O. Krejčí, S. Dimosthenous, S. R. Kachel, M. Chen, A. S. Foster, U. Koert, P. Liljeroth et al., Science, 2021, 372, 852–856.
  • Berber et al. 2004 S. Berber, E. Osawa and D. Tománek, Phys. Rev. B, 2004, 70, 085417.
  • Hou et al. 2022 L. Hou, X. Cui, B. Guan, S. Wang, R. Li, Y. Liu, D. Zhu and J. Zheng, Nature, 2022, 606, 507–510.
  • Paras et al. 2022 Paras, K. Yadav, P. Kumar, D. R. Teja, S. Chakraborty, M. Chakraborty, S. S. Mohapatra, A. Sahoo, M. M. Chou, C.-T. Liang et al., Nanomaterials, 2022, 13, 160.
  • Júnior et al. 2023 M. P. Júnior, W. F. da Cunha, W. F. Giozza, R. T. de Sousa Junior and L. R. Junior, FlatChem, 2023, 37, 100469.
  • Zhang and Guo 2024 Y.-F. Zhang and J. Guo, Int. J. Hydrogen Energy, 2024, 50, 1004–1014.
  • Tan et al. 2024 Y. Tan, X. Tao, Y. Ouyang and Q. Peng, Int. J. Hydrogen Energy, 2024, 50, 738–748.
  • Majidi 2024 R. Majidi, Mater. Today Commun., 2024, 38, 107641.
  • Li et al. 2023 J. Li, M. Han, S. Zhao, T. Li, T. Yu, Y. Zhang, H.-K. Tang and Q. Peng, J. Compos. Sci., 2023, 7, 490.
  • Majidi and Ayesh 2024 R. Majidi and A. I. Ayesh, Adv. Theory Simul., 2024, 2400131.
  • Tan et al. 2024 Y. Tan, X. Tao, Y. Ouyang and Q. Peng, Int. J. Hydrogen Energy, 2024, 50, 738–748.
  • Felix and Pereira 2018 I. M. Felix and L. F. C. Pereira, Sci. Rep., 2018, 8, 1–10.
  • Mortazavi et al. 2014 B. Mortazavi, M. Pötschke and G. Cuniberti, Nanoscale, 2014, 6, 3344–3352.
  • Mortazavi and Ahzi 2013 B. Mortazavi and S. Ahzi, Carbon, 2013, 63, 460–470.
  • Mortazavi et al. 2016 B. Mortazavi, O. Rahaman, T. Rabczuk and L. F. C. Pereira, Carbon, 2016, 106, 1–8.
  • Thompson et al. 2022 A. P. Thompson, H. M. Aktulga, R. Berger, D. S. Bolintineanu, W. M. Brown, P. S. Crozier, P. J. in’t Veld, A. Kohlmeyer, S. G. Moore, T. D. Nguyen et al., Comp. Phys. Comm., 2022, 271, 108171.
  • Brenner et al. 2002 D. W. Brenner, O. A. Shenderova, J. A. Harrison, S. J. Stuart, B. Ni and S. B. Sinnott, J. Phys.: Condens. Matter., 2002, 14, 783.
  • Ong and Pop 2011 Z.-Y. Ong and E. Pop, Phys. Rev. B, 2011, 84, 075471.
  • Mu et al. 2014 X. Mu, X. Wu, T. Zhang, D. B. Go and T. Luo, Sci. Rep., 2014, 4, 3909.
  • Barbarino et al. 2015 G. Barbarino, C. Melis and L. Colombo, Phys. Rev. B, 2015, 91, 035416.
  • Wei et al. 2022 N. Wei, Y. Chen, K. Cai, Y. Zhang, Q. Pei, J.-C. Zheng, Y.-W. Mai and J. Zhao, Green Energy Environ., 2022, 7, 86–94.
  • Gale 1997 J. D. Gale, J. Chem. Soc., Faraday Trans., 1997, 93, 629–637.
  • Gale and Rohl 2003 J. D. Gale and A. L. Rohl, Mol. Simul., 2003, 29, 291–341.
  • Tersoff 1988 J. Tersoff, Phys. Rev. B, 1988, 37, 6991.
  • Lindsay and Broido 2010 L. Lindsay and D. Broido, Phys. Rev. B, 2010, 81, 205441.
  • Müller-Plathe 1997 F. Müller-Plathe, J. Chem. Phys., 1997, 106, 6082–6085.
  • Tromer et al. 2020 R. M. Tromer, I. M. Felix, A. Freitas, S. Azevedo and L. F. C. Pereira, Comput. Mater. Sci., 2020, 172, 109338.
  • Tromer et al. 2020 R. M. Tromer, A. Freitas, I. M. Felix, B. Mortazavi, L. Machado, S. Azevedo and L. F. C. Pereira, Phys. Chem. Chem. Phys., 2020, 22, 21147–21157.
  • Pereira and Felix 2022 L. F. C. Pereira and I. M. Felix, J. Phys.: Conf. Ser., 2022, p. 012008.
  • Lindsay et al. 2010 L. Lindsay, D. Broido and N. Mingo, Phys. Rev. B, 2010, 82, 115427.
  • Xu et al. 2015 W. Xu, G. Zhang and B. Li, J. Chem. Phys., 2015, 143, 154703.
  • Pereira et al. 2016 L. F. C. Pereira, B. Mortazavi, M. Makaremi and T. Rabczuk, RSC Adv., 2016, 6, 57773–57779.
  • Taheri et al. 2021 A. Taheri, S. Pisana and C. V. Singh, Phys. Rev. B, 2021, 103, 235426.
  • Schelling et al. 2002 P. K. Schelling, S. R. Phillpot and P. Keblinski, Phys. Rev. B, 2002, 65, 144306.
  • Felix and Pereira 2020 I. M. Felix and L. F. C. Pereira, Carbon, 2020, 160, 335–341.
  • Felix and Pereira 2022 I. M. Felix and L. F. C. Pereira, Int. J. Heat Mass Transf., 2022, 186, 122464.
  • Bae et al. 2013 M.-H. Bae, Z. Li, Z. Aksamija, P. N. Martin, F. Xiong, Z.-Y. Ong, I. Knezevic and E. Pop, Nat. Commun., 2013, 4, 1734.
  • Balandin et al. 2008 A. A. Balandin, S. Ghosh, W. Bao, I. Calizo, D. Teweldebrhan, F. Miao and C. N. Lau, Nano Lett., 2008, 8, 902–907.
  • Ghosh et al. 2008 S. Ghosh, I. Calizo, D. Teweldebrhan, E. P. Pokatilov, D. L. Nika, A. A. Balandin, W. Bao, F. Miao and C. N. Lau, Appl. Phys. Lett., 2008, 92, 151911.
  • Faugeras et al. 2010 C. Faugeras, B. Faugeras, M. Orlita, M. Potemski, R. R. Nair and A. Geim, ACS nano, 2010, 4, 1889–1892.
  • Cai et al. 2010 W. Cai, A. L. Moore, Y. Zhu, X. Li, S. Chen, L. Shi and R. S. Ruoff, Nano Lett., 2010, 10, 1645–1651.
  • Lee et al. 2011 J.-U. Lee, D. Yoon, H. Kim, S. W. Lee and H. Cheong, Phys. Rev. B, 2011, 83, 081419.
  • Tromer et al. 2023 R. Tromer, I. Felix, L. Pereira, M. da Luz, L. R. Junior and D. Galvao, Phys. Chem. Chem. Phys., 2023, 25, 28703–28715.
  • Tromer et al. 2022 R. M. Tromer, L. F. C. Pereira, M. Ferreira and M. da Luz, J.Appl. Phys., 2022, 131, 115704.
  • Tabarraei 2015 A. Tabarraei, Comput. Mater. Sci., 2015, 108, 66–71.
  • Yuan et al. 2019 C. Yuan, J. Li, L. Lindsay, D. Cherns, J. W. Pomeroy, S. Liu, J. H. Edgar and M. Kuball, Commun. Phys., 2019, 2, 43.
  • He et al. 2020 J. He, Y. Ouyang, C. Yu, P. Jiang, W. Ren and J. Chen, Chinese Phys. B, 2020, 29, 126503.
  • Ying et al. 2022 P. Ying, T. Liang, Y. Du, J. Zhang, X. Zeng and Z. Zhong, Int. J. Heat Mass Transf., 2022, 183, 122060.
  • Mortazavi et al. 2023 B. Mortazavi, F. Shojaei and X. Zhuang, Mater. Today Nano, 2023, 21, 100280.
  • Peimyoo et al. 2015 N. Peimyoo, J. Shang, W. Yang, Y. Wang, C. Cong and T. Yu, Nano Res., 2015, 8, 1210–1221.
  • Yan et al. 2014 R. Yan, J. R. Simpson, S. Bertolazzi, J. Brivio, M. Watson, X. Wu, A. Kis, T. Luo, A. R. Hight Walker and H. G. Xing, ACS Nano, 2014, 8, 986–993.
  • Kuang et al. 2016 Y. D. Kuang, L. Lindsay, S. Q. Shi and G. Zheng, Nanoscale, 2016, 8, 3760–3767.
  • Qin et al. 2015 G. Qin, Q.-B. Yan, Z. Qin, S.-Y. Yue, M. Hu and G. Su, Phys. Chem. Chem. Phys., 2015, 17, 4854–4858.
  • Yin et al. 2021 S. Yin, W. Zhang, C. Tan, L. Chen, J. Chen, G. Li, H. Zhang, Y. Zhang, W. Wang and L. Li, J. Phys. Chem. C, 2021, 125, 16129–16135.
  • Seol et al. 2010 J. H. Seol, I. Jo, A. L. Moore, L. Lindsay, Z. H. Aitken, M. T. Pettes, X. Li, Z. Yao, R. Huang, D. Broido et al., Science, 2010, 328, 213–216.
  • Pereira and Donadio 2013 L. F. C. Pereira and D. Donadio, Phys. Rev. B, 2013, 87, 125424.
  • Ziman 2001 J. Ziman, Electrons and Phonons: The Theory of Transport Phenomena in Solids, Oxford University Press, 2001.